Calculate the amount and compound interest on
₹ 10000 for 1 year at 8% per annum compounded half yearly. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Principal (P) =₹10,000
Rate = 8% per annum or 4% per half year Number of years = 1 year
There are 2 half years in 1 year.
A =P (1 +R/100 )^n
=₹ [10000(1+ 4/100 )^2]
=₹ [10000 × (1/25)^2]
=₹ [10000 (26/25)^2]
=₹ 10816
C.I. = A − P = ₹ 10816 − ₹10000 =₹816

​​Thanks